# Topological Center of the Double Dual of the Orlicz Figà-Talamanca Herz algebra

### v1

@@ 5. Topological center
Since the restriction map is a bounded projection which is also a multiplier, the result is a direct consequence of [41, Theorem 4.2]. ∎
Here is the main result of this article concerning the topological center of the double dual of This theorem serves as an Orlicz analogue of Miao’s result on algebras [41, Theorem 4.4].
@@
Let be an amenable group and satisfies the MA condition. Then if and only if for any open compact subgroup of
@@
So, Now, by Theorem 4.5, the restriction of (say ) to is in i.e., for some where is the canonical map from to Since is amenable, by [32, Theorem 3.1], the algebra has a bounded approximate identity, say and further, by Theorem 3.4, the algebra has a right identity, say Moreover, it is clear from the proof of [4, Proposition III.28.7] that is the cluster point of Thus, for any since the function is in we have,
@@
In the subsequent results, we study the topological center of the Banach algebra From [7, Corollary 3.4], we recall that the space is continuously embedded in under the assumption that the group is amenable. Further, it is clear from the identification in [7, Theorem 3.2 and Corollary 3.4] that for any the element in is given by
@@
To prove this lemma, it is enough to show that the space is contained in the center of as in [35, Proposition 4.5]. Now, let and Then, for any and we have,
@@
The final theorem of the article deals with the topological center of the algebra and is analogous to Theorem 5.2. The proof of the theorem follows verbatim to [41, Corollary 4.6 (ii)] and hence, we omit it. The main ingredients of the proof are Lemma 5.3 and the following two facts:
@@
For an amenable group the space which is the predual of is contained in This follows from [31, Corollary 4.13 (i)], since is amenable.
@@
Let be an amenable group. Then if and only if for any open compact subgroup of

### v2

@@ 5. Topological center
Since the restriction map is a bounded projection which is also a multiplier, the result is a direct consequence of [44, Theorem 4.2]. ∎
Here is the main result of this article concerning the topological center of the double dual of This theorem serves as an Orlicz analogue of Miao’s result on algebras [44, Theorem 4.4].
@@
Let be an amenable group and suppose that satisfies the MA condition. Then if and only if for every open compact subgroup of
@@
So, Now, by Theorem 4.5, the restriction of (say ) to is in i.e., for some where is the canonical map from to Since is amenable, by [33, Theorem 3.1], the algebra has a bounded approximate identity, say and further, by Theorem 3.4, the algebra has a right identity, say Moreover, it is clear from the proof of [5, Proposition III.28.7] that is the cluster point of Thus, for any since the function is in we have,
@@
In the subsequent results, we study the topological center of the Banach algebra From [8, Corollary 3.4], we recall that the space is continuously embedded in under the assumption that the group is amenable. Further, it is clear from the identification in [8, Theorem 3.2 and Corollary 3.4] that for any the element in is given by
@@
To prove this lemma, it is enough to show that the space is contained in the center of as in [36, Proposition 4.5]. Now, let and Then, for any and we have,
@@
The final theorem of the article deals with the topological center of the algebra and is analogous to Theorem 5.2. The proof of the theorem follows verbatim to [44, Corollary 4.6 (ii)] and hence, we omit it. The main ingredients of the proof are Lemma 5.3 and the following two facts:
@@
For an amenable group the space which is the predual of is contained in This follows from [32, Corollary 4.13 (i)], since is amenable.
@@
Let be an amenable group. Then if and only if for every open compact subgroup of